# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v1

@@ 4.3 A strong RL baseline: PPO, and why it falls short in our benchmark
Why include PPO, and how we match the setting. Proximal Policy Optimization (PPO) is a widely used and robust model-free policy-gradient baseline for continuous control (schulman2017proximal). We include PPO to answer a concrete question: can a generic, well-tuned model-free RL method recover the decision-time $q$-optimal $\theta$-blind allocation in our high-dimensional drift-uncertainty benchmark under comparable simulation budgets? This comparison is especially informative in our static Gaussian benchmark because the target decision-time rule is structurally simple (constant and available in closed form), so performance gaps primarily reflect optimization difficulty and credit assignment rather than policy-class expressiveness.
Since classical HJB solvers and value-function-based deep PDE surrogates are not practical baselines in the high-dimensional uncertain regime targeted here (Section 2.4), PPO serves as a strong simulation-only comparator that operates on the same sampled trajectories without exploiting value-function PDE structure. For a fair comparison, PPO is trained on the same Euler simulator and time discretization as our PG–DPO pipeline, under the same deployability restriction (the policy never observes the latent $\theta$), and under the same terminal-utility objective. We also use the same action cap $u_{\max}$ (with the same dimension-scaling convention) so that exploration ranges are comparable across $d$. Implementation details are deferred to the appendix and code release.
@@
Why PPO underperforms in this benchmark. The gap is not evidence that PPO is intrinsically weak; rather, it reflects that our benchmark stresses regimes where a generic likelihood-ratio policy gradient is statistically disadvantaged compared to pathwise/adjoint-based updates. With terminal utility as the only reward, PPO faces a long-horizon credit-assignment problem whose gradient variance grows with both horizon and action dimension. Sampling $\theta\sim q$ further creates episode-wise heterogeneity under a single $\theta$-blind policy, inducing additional variance and potential cancellation across parameter draws. In contrast, Stage 1 exploits backpropagation through the differentiable simulator (pathwise gradients), and Stage 2 leverages the affine-in-control Pontryagin structure through a $q$-aggregated projection, replacing a noisy high-dimensional policy-gradient update by a structured stationarity correction that is tailored to the $\theta$-blind ex–ante objective.
Under matched simulation budgets in our latent-$\theta$, $\theta$-blind benchmark, a generic model-free PPO baseline does not reliably recover the decision-time $q$-optimal allocation (Table LABEL:tab:crra_stage12_distill_ppo_landscape), motivating structure-exploiting alternatives—pathwise gradients, costates, and the $q$-aggregated Pontryagin projection—as in PG–DPO and P–PGDPO.
@@ 5 Recovering Intertemporal Hedging Demand in Factor-Driven Markets
Sections 4 stressed scaling under static drift uncertainty, where the target $q$-reference is time-homogeneous and largely myopic. Here we shift the focus to an economic target: recovering the intertemporal hedging demand induced by factor-driven investment opportunities when return shocks are correlated with factor shocks (campbell2002strategic; xia2001learning).
We use the mean-reverting Gaussian premium benchmark of Section 2.3.2. Decision-time statistical uncertainty enters through the (uncertain) initial premium state $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$, while a nonzero return–factor correlation $\rho$ generates hedging demand through the cross term $M_{\mathrm{cross}}(T)$ in (51). Crucially, we enforce a deployable restriction aligned with Section 3: the policy is $Y$-blind and does not observe the realized $Y_{0}$ nor the path $(Y_{t})$.
We compare: (i) Stage 1 PG–DPO (deployable end-to-end learning; Section 3.1), (ii) Stage 2 $q$-aggregated Pontryagin projection (post-hoc correction in residual form; Sections 3.2 and 3.3.1), (iii) interactive distillation (amortized projection guidance; Section 3.3.2), and (iv) a model-free PPO baseline trained under the same deployable $Y$-blind observation restriction. Performance is measured by decision-time RMSE against the analytic constant-portfolio OU reference (51) (which reduces to the independence-case benchmark when $\rho=0$).
In addition to the full allocation error, this benchmark provides a natural myopic + hedging decomposition (driven by return–factor correlation). We therefore report (a) the RMSE of the full decision-time allocation for all methods (including PPO), and (b) component-wise diagnostics for the projected (Stage 2) rules: RMSE of the hedging component (Table 1) and, in the appendix, the RMSE of the myopic component (Table 2) and the cosine similarity of the hedging direction (Table 3). Since PPO does not expose a compatible myopic/hedging decomposition for these diagnostics, we include it only in the full RMSE table.
To keep the main text focused, we include the full RMSE table (Table LABEL:tab:ou_rmse_s0_sweep_landscape) and the hedging-RMSE table (Table 1) in Section 5.2; the remaining two diagnostic tables are deferred to the appendix (Section F).
@@ 5.1 Experimental setting
$Y$-blind deployability (and what uses $Y$). Throughout this section, all reported policies are deployable and $Y$-blind: the control is a function of observable wealth and time-to-go only, and never takes the realized initial premium $Y_{0}$ nor the factor path $(Y_{t})$ as an input (including the PPO baseline). The latent premium factor is sampled and propagated only inside the simulator to generate trajectories and to form Monte Carlo averages used by the stage 2 projection (and by the teacher in distillation). Any $Y$-indexed quantities are used only for offline evaluation and diagnostics.
OU premium market with a hedging channel. We adopt the OU premium benchmark of Section 2.3.2. Let $Y_{t}\in\mathbb{R}^{m}$ be a mean-reverting premium factor and $R_{t}\in\mathbb{R}^{d}$ the risky excess returns:
$dY_{t}$ $=\kappa(\bar{y}-Y_{t})\,dt+\Xi\,dW_{t}^{Y},\hskip 18.49988ptY_{0}\sim\mathcal{N}(m_{0},P_{0}),$
@@
A nonzero $\rho$ induces intertemporal hedging demand and enters the CRRA decision-time reference through the cross-covariance term $M_{\mathrm{cross}}(T)$ in (51). When $\rho=0$ (independent return and factor shocks), the hedging channel vanishes ($M_{\mathrm{cross}}(T)=0$) and the reference reduces to the independence-case benchmark.
Decision-time uncertainty geometry for $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$. We control the magnitude of decision-time statistical uncertainty by a scalar $s_{0}>0$ and construct $P_{0}$ from an identification-motivated baseline
@@
Two-stage solver, amortization, and evaluation protocol. We use the two-stage pipeline of Section 3. Stage 1 trains a deployable policy by stochastic gradient ascent using pathwise/BPTT gradients (Section 3.1). Stage 2 applies the $q$-aggregated Pontryagin projection computed under a warm-up policy (Section 3.2), implemented in the residual/control-variate form (Section 3.3.1). Interactive distillation amortizes the projected teacher into a fast deployable policy network (Section 3.3.2). As a model-free baseline, we also train a PPO policy under the same $Y$-blind observation restriction and report its decision-time full RMSE in Table LABEL:tab:ou_rmse_s0_sweep_landscape.
We sweep $d\in\{5,10,50,100\}$ (one fixed market instance per $d$), train for $5000$ epochs, and evaluate every $100$ epochs. Unless stated otherwise we set $\gamma=2$, $r=0.03$, $\kappa=1.0$, $\xi_{\mathrm{scale}}=0.25$, and $\rho=0.5$. We evaluate the decision-time action at $t=0$ and report RMSE to the analytic constant-portfolio OU reference (51).
In addition to the full allocation error, we use the natural myopic + hedging decomposition induced by the OU factor structure. We report the RMSE of the full decision-time allocation for all methods, and component-wise diagnostics for the projected (Stage 2) rules, including the RMSE of the hedging component and (in the appendix) the RMSE of the myopic component and cosine similarity of the hedging direction. To reduce noise from stochastic optimization, for each condition we summarize each metric by a tail median over the last six evaluation checkpoints.
@@ 5.2 Results: hedging-demand recovery, amortization, and robustness to decision-time uncertainty
We report decision-time RMSE at $t=0$ against the analytic OU reference (51). To reduce noise from stochastic optimization, we summarize each condition by a tail median over the last six evaluation checkpoints. Table LABEL:tab:ou_rmse_s0_sweep_landscape reports the full decision-time RMSE for all deployable objects (Stage 1 and Stage 1+Stage 2, with and without distillation), and also includes a model-free PPO baseline trained under the same $Y$-blind deployability restriction. To isolate the economic channel of interest, Table 1 reports the RMSE of the hedging component for the post-hoc projected (Stage 2) rules. Two additional diagnostics—the myopic-component RMSE and the hedging-direction cosine similarity—are deferred to the appendix (Tables 2 and 3). Since PPO does not expose a compatible myopic/hedging decomposition in our diagnostic protocol, we report it only in the full-RMSE table.
Amortization, robustness, and the PPO baseline. Interactive distillation improves the deployable Stage 1 policy relative to the basic PG–DPO run, while the most accurate object remains the post-hoc projected policy (Table LABEL:tab:ou_rmse_s0_sweep_landscape). This matches the intended division of labor in Section 3.3: Stage 2 provides a structured stationarity-correction signal through the aggregated Pontryagin projection, and distillation amortizes that correction into a single forward pass, up to policy-class approximation limits.
Finally, the PPO baseline remains far from the analytic OU reference under the same $Y$-blind deployability restriction, with degradation that becomes especially pronounced at larger $d$ (Table LABEL:tab:ou_rmse_s0_sweep_landscape). This is consistent with PPO facing a terminal-only credit-assignment problem under latent-factor heterogeneity, in contrast to the pathwise-sensitivity and affine-in-control correction exploited by our two-stage pipeline. Since PPO does not provide a compatible myopic/hedging decomposition under our evaluation protocol, we include it only in the full-RMSE table.
@@
Table 1: Decision-time RMSE at $t=0$ for the hedging component of the OU decision-time reference, evaluated on the post-hoc projected (Stage 2) policies (tail median over the last six evaluation checkpoints). Component-wise diagnostics are reported for Stage 2 since Stage 1 does not explicitly output a myopic/hedging decomposition.
In a factor-driven market where return–factor correlation induces intertemporal hedging demand, the proposed two-stage pipeline recovers the analytic OU decision-time reference with high accuracy: projection provides the dominant gains, and distillation improves deployable policies by amortizing projection guidance. In contrast, a model-free PPO baseline does not reliably match the analytic reference in this $Y$-blind setting.
@@ 6 Conclusion
Methodologically, we extended Pontryagin–Guided Direct Policy Optimization (PG–DPO) to the latent-parameter setting by sampling $\theta$ only inside the simulator and computing exact discrete-time gradients via BPTT (Section 3.1), and we leveraged the BPTT–PMP correspondence to extract the costate objects needed for structured control updates (Theorem 2). Building on the $q$-aggregated stationarity, we proposed uncertainty-aware projected PG–DPO (P–PGDPO), which aggregates Monte Carlo Pontryagin quantities across $\theta\sim q$ and projects them onto the deployable first-order condition to obtain a single $\theta$-blind rule (Section 3.2). We established a residual-based ex–ante policy-gap bound under local stability of the aggregated projection map, with discretization and Monte Carlo errors made explicit (Theorem 3). In experiments with finite-sample uncertainty, projection improves stability and accuracy in high dimensions and exhibits a two-time-scale stabilization effect (costates versus policies), while interactive distillation amortizes the projection into a fast deployable network (Sections 4 and 5; Section 3.3).
Several extensions are natural. A first direction is to allow time-varying uncertainty descriptions $q_{t}$ (e.g., produced by an external filter) and connect the present fixed-$q$ projection to belief-aware decision rules (Remark 2, Appendix A). A second direction is to incorporate realistic frictions and constraints (transaction costs, leverage and short-sale limits) and develop certified or regularized projection steps when mixed-moment estimation becomes fragile (Section 2.4; Section 4.2; Appendix D.5). Finally, applying the framework to large cross-sectional datasets with modern estimation pipelines would further clarify the practical benefits of inference-agnostic, simulation-only optimization under parameter uncertainty (Section 1).
@@ Appendix A Online uncertainty updates: Kalman–Bucy filtering and a plug-in decision-time benchmark
Purpose and scope. Sections 2.3.1 and 2.3.2 focus on decision-time benchmarks in which an uncertainty description $q$ is treated as given and the investor optimizes under the corresponding $\theta$-blind deployability constraint. In practice, however, new data arrive and the uncertainty description is updated over time by an external estimation/filtering engine, a viewpoint that aligns with learning/estimation-risk portfolio choice and Bayesian decision-time formulations (barberis2000investor; pastor2000portfolio; xia2001learning). This subsection records a simple linear–Gaussian example in which such an updated description $q_{t}$ arises endogenously via a Kalman–Bucy filter (a canonical partially observed diffusion setting; see, e.g., bensoussan1985optimal; pham2017dynamic), and then formalizes a plug-in workflow: at each decision time, treat the current uncertainty description $q_{t}$ as given and compute a decision-time optimal control under that $q_{t}$. We emphasize that solving the fully optimal partial-observation (belief-state) control problem is not the goal of this paper; rather, we view the resulting $q_{t}$ as an external input to decision-time optimization. In particular, our simulation-based Pontryagin-guided solvers developed later (Section 3) can be used as inner-loop engines that are refreshed whenever a new uncertainty description $q_{t}$ becomes available.
@@
Thus, even though the posterior $q_{t}$ is a distribution-valued object, in this affine/Gaussian regime it is fully characterized by the finite-dimensional sufficient statistics $(\hat{Y}_{t},P_{t})$, with $P_{t}$ evolving deterministically via (90); this is the prototypical setting in which belief-state control reduces to finite-dimensional sufficient statistics (bensoussan1985optimal; pham2017dynamic).
@@
One may interpret (96) as a receding-horizon decision-time policy driven by an externally updated uncertainty description, consistent with the general “update beliefs, then optimize” workflow used in Bayesian/learning-based portfolio choice (barberis2000investor; pastor2000portfolio; xia2001learning).

### v2

@@ 4.3 A strong RL baseline: PPO, and why it falls short in our benchmark
Why include PPO, and how we match the setting. Proximal Policy Optimization (PPO) is a widely used and robust model-free policy-gradient baseline for continuous control (Schulman et al., 2017). We include PPO to answer a concrete question: can a generic, well-tuned model-free RL method recover the decision-time $q$-optimal $\theta$-blind allocation in our high-dimensional drift-uncertainty benchmark under comparable simulation budgets? This comparison is especially informative in our static Gaussian benchmark because the target decision-time rule is structurally simple (constant and available in closed form), so performance gaps primarily reflect optimization difficulty and credit assignment rather than policy-class expressiveness.
Since classical HJB solvers and value-function-based deep PDE surrogates are not practical baselines in the high-dimensional uncertain regime targeted here (Section B), PPO serves as a strong simulation-only comparator that operates on the same sampled trajectories without exploiting value-function PDE structure. For a fair comparison, PPO is trained on the same Euler simulator and time discretization as our PG–DPO pipeline, under the same deployability restriction (the policy never observes the latent $\theta$), and under the same terminal-utility objective. We match simulation budgets at the level of total environment interactions (simulated trajectories $\times$ time steps); implementation and PPO tuning details are deferred to the appendix and code release.
@@
Why PPO underperforms in this benchmark. The gap is not evidence that PPO is intrinsically weak; rather, it reflects that our benchmark stresses regimes where a generic likelihood-ratio policy gradient is statistically disadvantaged compared to pathwise/adjoint-based updates.
Terminal-reward credit assignment. With terminal utility as the only reward signal, PPO faces a long-horizon credit-assignment problem: most intermediate transitions carry zero reward, so advantage estimates inherit high variance and weak learning signals, especially in higher dimensions.
One contributor is terminal-reward credit assignment: with terminal utility as the only reward signal, most intermediate transitions carry no direct learning signal, so advantage estimates become noisy and the gradient signal weakens as the horizon and dimension grow. A second contributor is high-dimensional action noise and variance scaling: in continuous control, exploration is typically implemented through stochastic policies (e.g., diagonal Gaussians), and as $d$ increases the effective exploration noise and likelihood-ratio gradient variance can inflate rapidly, making it difficult to consistently recover the precise constant-fraction structure of the $q$-optimal $\theta$-blind rule under fixed simulation budgets.
In contrast, Stage 1 exploits backpropagation through the differentiable simulator (pathwise gradients), and Stage 2 leverages the affine-in-control Pontryagin structure through a $q$-aggregated projection, replacing a noisy high-dimensional policy-gradient update by a structured stationarity correction that is tailored to the $\theta$-blind ex–ante objective.
@@ 5 Recovering Intertemporal Hedging Demand in Factor-Driven Markets
Sections 4 stressed scaling under static drift uncertainty, where the target $q$-reference is time-homogeneous and largely myopic. Here we shift the focus to an economic target: recovering the intertemporal hedging demand induced by factor-driven investment opportunities when return shocks are correlated with factor shocks (Campbell and Viceira, 2002; Xia, 2001).
We adopt the mean-reverting Gaussian premium benchmark of Section 2.3.2. At decision time ($t=0$), the investor is assumed to have access only to an interval/posterior uncertainty description $q(dy)$ for the initial premium state $Y_{0}$ (e.g., $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$), and we restrict attention to deployable $Y$-blind policies that depend only on observable wealth and time-to-go. When return and factor shocks are correlated ($\rho\neq 0$), the analytic OU decision-time reference (43) contains the cross term $M_{\mathrm{cross}}(\tau)$ (with $\tau:=T-t$) that encodes intertemporal hedging demand.
We evaluate the two-stage PG–DPO / $q$-aggregated Pontryagin projection pipeline (and interactive distillation) against a model-free PPO baseline, all under the same $Y$-blind deployability restriction. Performance is measured by decision-time RMSE at $t=0$ relative to the analytic constant-portfolio OU reference (43).
@@ 5.1 Experimental setting
Decision-time interval/posterior estimate for $Y_{0}$ and deployability. Throughout this section, all reported policies are deployable and $Y$-blind: the control is a function of observable wealth and time-to-go only, and never takes as input (i) a point estimate of $Y_{0}$, (ii) the realized initial premium $Y_{0}$, or (iii) the factor path $(Y_{t})$ (including the PPO baseline). At deployment, the investor is assumed to have access only to an uncertainty description $q(dy)$ for $Y_{0}$ (interval/posterior output of an external estimation pipeline), and commits to a single policy optimized ex ante under $Y_{0}\sim q$ (without filtering/updating over $[0,T]$).
The latent premium factor is sampled and propagated only inside the simulator to generate trajectories and to form Monte Carlo averages used by the stage 2 projection (and by the teacher in distillation). Any $Y$-indexed quantities are used only for offline evaluation and diagnostics.
OU premium market with a hedging channel. We adopt the OU premium benchmark of Section 2.3.2. To avoid a notational clash with other uses of $\kappa$ elsewhere in the paper, we denote the OU mean-reversion rate by $\kappa_{Y}$ in this section. Let $Y_{t}\in\mathbb{R}^{m}$ be a mean-reverting premium factor and $R_{t}\in\mathbb{R}^{d}$ the risky excess returns:
$dY_{t}$ $=\kappa_{Y}(\bar{y}-Y_{t})\,dt+\Xi\,dW_{t}^{Y},\hskip 18.49988ptY_{0}\sim\mathcal{N}(m_{0},P_{0}),$
@@
Here $\rho\in\mathbb{R}^{d\times m}$ denotes the instantaneous return–factor shock correlation matrix, satisfying the feasibility condition $\rho^{\top}\rho\preceq I_{m}$. A nonzero $\rho$ induces intertemporal hedging demand and enters the CRRA decision-time reference through the cross-covariance term $M_{\mathrm{cross}}(\tau)$ in (43), where $\tau:=T-t$ is time-to-go (so at decision time $t=0$, $\tau=T$). When $\rho=0$ (independent return and factor shocks), the hedging channel vanishes ($M_{\mathrm{cross}}(\tau)=0$) and the reference reduces to the independence-case benchmark.
Decision-time uncertainty geometry for $Y_{0}\sim\mathcal{N}(m_{0},P_{0})$. We control the magnitude of decision-time estimation uncertainty by a scalar $s_{0}>0$ and construct $P_{0}$ from an identification-motivated baseline
@@
Two-stage solver, amortization, and evaluation protocol. We use the two-stage pipeline of Section 3. Stage 1 trains a deployable policy by stochastic gradient ascent using pathwise/BPTT gradients (Section 3.1). Stage 2 applies the $q$-aggregated Pontryagin projection computed under a warm-up policy (Section 3.2), implemented in the residual/control-variate form (Section F.1). Interactive distillation amortizes the projected teacher into a fast deployable policy network (Section F.2). As a model-free baseline, we also train a PPO policy under the same $Y$-blind observation restriction and report its decision-time full RMSE in Table LABEL:tab:ou_rmse_s0_sweep_landscape.
We sweep $d\in\{5,10,50,100\}$ (one fixed market instance per $d$, following the same market-generation protocol used in Section 4), train for $5000$ epochs, and evaluate every $100$ epochs. We evaluate the decision-time action at $t=0$ (i.e., $\tau=T$) and report RMSE to the analytic constant-portfolio OU reference (43).
Unless stated otherwise we set $\gamma=2$, $r=0.03$, $\kappa_{Y}=1.0$, $\xi_{\mathrm{scale}}=0.25$, and a correlation magnitude $\rho_{0}=0.5$. We instantiate the model parameters as follows: (i) we take $\Xi=\xi_{\mathrm{scale}}I_{m}$, (ii) we implement the return–factor correlation by choosing a matrix $\rho=\rho_{0}Q$ with $Q\in\mathbb{R}^{d\times m}$ having orthonormal columns ($Q^{\top}Q=I_{m}$), ensuring $\rho^{\top}\rho=\rho_{0}^{2}I_{m}\preceq I_{m}$. (A simple choice is $Q=[I_{m};0]\in\mathbb{R}^{d\times m}$ when $m\leq d$; any fixed orthonormal-column $Q$ yields an equivalent correlation magnitude.) We fix the horizon to $T=1.5$ in this section.
In addition to the full allocation error, we use the myopic+hedging decomposition induced by the OU structure, and report component-wise diagnostics for the projected (Stage 2) rules. To reduce noise from stochastic optimization, for each condition we summarize each metric by a tail median over the last six evaluation checkpoints.
@@ 5.2 Results: hedging-demand recovery, amortization, and robustness to decision-time uncertainty
Metrics and diagnostics. Table LABEL:tab:ou_rmse_s0_sweep_landscape reports the full decision-time RMSE at $t=0$ for all deployable policies, including the PPO baseline. To isolate the economic hedging channel, we use the OU-induced decomposition of the analytic reference
$\pi^{\mathrm{ref}}(\tau)=\pi^{\mathrm{myo}}(\tau)+\pi^{\mathrm{hedge}}(\tau),$
where $\pi^{\mathrm{myo}}$ is the independence-case reference obtained by setting $\rho=0$ (equivalently, dropping $M_{\mathrm{cross}}(\tau)$ in (43)), and $\pi^{\mathrm{hedge}}:=\pi^{\mathrm{ref}}-\pi^{\mathrm{myo}}$ is the residual $\rho$-dependent term. We report the hedging-component RMSE for the projected (Stage 2) rules in Table 1; the myopic-component RMSE and hedging-direction cosine similarity are deferred to Appendix H. Since Stage 1 and PPO do not expose a compatible myopic/hedging split under our diagnostic protocol, component-wise diagnostics are reported only for Stage 2.
@@
Table 1: Decision-time RMSE at $t=0$ for the hedging component of the OU decision-time reference, evaluated on the post-hoc projected (Stage 2) policies (tail median over the last six evaluation checkpoints). We use the decomposition $\pi^{\mathrm{ref}}(\tau)=\pi^{\mathrm{myo}}(\tau)+\pi^{\mathrm{hedge}}(\tau)$ where $\pi^{\mathrm{myo}}$ is the $\rho=0$ reference and $\pi^{\mathrm{hedge}}$ is the residual $\rho$-dependent term (cf. (43)). Component-wise diagnostics are reported for Stage 2 since Stage 1 and PPO do not explicitly output a compatible myopic/hedging split under our protocol.
Amortization, robustness, and the PPO baseline. Interactive distillation improves the deployable Stage 1 policy relative to the basic PG–DPO run, while the most accurate object remains the post-hoc projected policy (Table LABEL:tab:ou_rmse_s0_sweep_landscape). This matches the intended division of labor in Section F: Stage 2 provides a structured stationarity-correction signal through the aggregated Pontryagin projection, and distillation amortizes that correction into a single forward pass, up to policy-class approximation limits.
Finally, the PPO baseline remains far from the analytic OU reference under the same $Y$-blind deployability restriction, and can be unstable in this terminal-only setting (Table LABEL:tab:ou_rmse_s0_sweep_landscape). This is consistent with PPO facing a difficult credit-assignment problem under latent-factor heterogeneity, in contrast to the pathwise-sensitivity and affine-in-control correction exploited by our two-stage pipeline. Since PPO does not provide a compatible myopic/hedging decomposition under our evaluation protocol, we include it only in the full-RMSE table.
@@ 6 Conclusion
Methodologically, we extended Pontryagin–Guided Direct Policy Optimization (PG–DPO) to the latent-parameter setting by sampling $\theta$ only inside the simulator and computing exact discrete-time gradients via BPTT (Section 3.1), and we leveraged the BPTT–PMP correspondence to extract the costate objects needed for structured control updates (Theorem 2). Building on the $q$-aggregated stationarity, we proposed uncertainty-aware projected PG–DPO (P–PGDPO), which aggregates Monte Carlo Pontryagin quantities across $\theta\sim q$ and projects them onto the deployable first-order condition to obtain a single $\theta$-blind rule (Section 3.2). We established a residual-based ex–ante policy-gap bound under local stability of the aggregated projection map, with discretization and Monte Carlo errors made explicit (Theorem 3). In experiments, projection provides the dominant gains—improving stability and decision-time accuracy in high-dimensional drift-uncertainty scaling tests and recovering intertemporal hedging demand in factor-driven markets—while interactive distillation amortizes projection into a fast deployable network (Sections 4 and 5; Section F).
Several extensions are natural. A first direction is to allow time-varying uncertainty descriptions $q_{t}$ (e.g., produced by an external filter) and connect the present fixed-$q$ projection to belief-aware decision rules (Remark 2, Appendix A). A second direction is to incorporate realistic frictions and constraints (transaction costs, leverage and short-sale limits) and develop certified or regularized projection steps when mixed-moment estimation becomes fragile (Section B; Section 4.2; Appendix E.5). Finally, applying the framework to large cross-sectional datasets with modern estimation pipelines would further clarify the practical benefits of inference-agnostic, simulation-only optimization under parameter uncertainty (Section 1).
@@ Appendix A Online uncertainty updates: Kalman–Bucy filtering and a plug-in decision-time benchmark
Purpose and scope. Sections 2.3.1 and 2.3.2 focus on decision-time benchmarks in which an uncertainty description $q$ is treated as given and the investor optimizes under the corresponding $\theta$-blind deployability constraint. In practice, however, new data arrive and the uncertainty description is updated over time by an external estimation/filtering engine, a viewpoint that aligns with learning/estimation-risk portfolio choice and Bayesian decision-time formulations (Barberis, 2000; Pástor, 2000; Xia, 2001). This subsection records a simple linear–Gaussian example in which such an updated description $q_{t}$ arises endogenously via a Kalman–Bucy filter (a canonical partially observed diffusion setting; see, e.g., Bensoussan and van Schuppen (1985); Pham and Wei (2017)), and then formalizes a plug-in workflow: at each decision time, treat the current uncertainty description $q_{t}$ as given and compute a decision-time optimal control under that $q_{t}$. We emphasize that solving the fully optimal partial-observation (belief-state) control problem is not the goal of this paper; rather, we view the resulting $q_{t}$ as an external input to decision-time optimization. In particular, our simulation-based Pontryagin-guided solvers developed later (Section 3) can be used as inner-loop engines that are refreshed whenever a new uncertainty description $q_{t}$ becomes available.
@@
Thus, even though the posterior $q_{t}$ is a distribution-valued object, in this affine/Gaussian regime it is fully characterized by the finite-dimensional sufficient statistics $(\hat{Y}_{t},P_{t})$, with $P_{t}$ evolving deterministically via (79); this is the prototypical setting in which belief-state control reduces to finite-dimensional sufficient statistics (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
@@
One may interpret (85) as a receding-horizon decision-time policy driven by an externally updated uncertainty description, consistent with the general “update beliefs, then optimize” workflow used in Bayesian/learning-based portfolio choice (Barberis, 2000; Pástor, 2000; Xia, 2001).